Use the simplex algorithm to find all optimal solutions to the following LP.

max z=2x1+x2
s.t. 4x1 + 2x2 ≤ 4

−2x1 + x2 ≤ 2

x1 ≥1

x1,x2 ≥0
Answer #1

Solution:

The problem is converted to canonical form by adding slack, surplus and artificial variables as appropiate 1. As the constraiEntering = x1, Departing = 41, Key Element = 1 + Rz(new) = R3(old) + R, (new) = R (old) - 4R3(new) + R (new) = Rz(old) + 2R;(+ Rz(new) = Rz(old) + R (new) Iteration-3 C2 0 XB MinRatio 0 0.25 0.5 0.5 0.25 z = 2 1 0.5 Z;-C; 0 0 0.5 Since all Z;-C;20 He